# Ringdown in Vaidya spacetimes: time-dependent frequencies, Penrose limit and time-domain analyses

### v2

@@ A. Penrose limit
$\omega^{\rm Re}_{\rm PL,ad}$ $:=$ $\frac{\Omega_{\rm ad}}{2}\left(2\ell+1\right),$ (131)
$\omega^{\rm Im}_{\rm PL,ad}$ $:=$ $\frac{\lambda_{\rm ad}}{2},$ (132)
$\mathcal{R}^{\rm ad}_{\rm PL}$ $:=$ $\left|\frac{\omega^{\rm Im}_{\rm PL,ad}}{\omega^{\rm Re}_{\rm PL,ad}}\right|.$ (133)
@@
Figure 6: The values of $\lambda_{\rm ad}$ and $\Omega_{\rm ad}$ with $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$ and $\mathcal{V}_{2}=100M_{1}$ as a function of $\mathcal{V}$ (left panel) and $\mathcal{V}_{\rm obs}$ (right panel). The null dust is accreting in the shaded region. The ratio $\lambda_{\rm ad}/\Omega_{\rm ad}$ is also plotted with the scale of the right-hand vertical axis as a reference.
We can find that the ratio $\lambda_{\rm ad}/\Omega_{\rm ad}$ is smaller than 1 in the accreting domain. Therefore, if the feature is reflected to the ringdown waveform, we may observe a temporary reduction of the value of $\mathcal{R}_{n}(2\ell+1)$ from the value of the Schwarzschild case. It would be interesting to note that we trivially obtain the Schwarzschild metric, taking the adiabatic approximation for the original Vaidya metric, since we ignore the time-dependence of the mass function. However, the adiabatic limit of the Penrose limit geometry reflects the time-dependence of the original geometry.
@@ B. Ringdown in Vaidya spacetimes and comparison with the Penrose limit analysis
We plot the values of $\Omega_{n}:=2\omega^{\rm Re}_{n}/(2\ell+1)$, $\lambda_{n}:=2\omega^{\rm Im}_{n}$ and $\mathcal{R}_{n}$, with $\mathcal{V}_{\rm obs}(\mathcal{V}_{n})$ being the horizontal axis value in Fig. 8. In the top panels, we can find that the values of $\Omega_{n}$ and $\lambda_{n}$ significantly deviate from $\Omega_{\rm ad}$ and $\lambda_{\rm ad}$, respectively. Since the qualitative behaviors are similar in both $\Omega_{n}$ and $\lambda_{n}$, these deviations mainly originate from the redshift effect, which equally causes the time delay to the real and imaginary parts of the frequency during the propagation from the UCOP to the observer. To extract the effects independent of the redshift, let us focus on the ratio $\mathcal{R}_{n}$ between the imaginary and real frequencies shown in the bottom panel of Fig. 8. For the bottom panel, since the predicted value from the PL analysis does not match the Schwarzschild case in both of the early and late times, we multiply the correction factor $\mathcal{C}$ (blue dashed line), so that it overlaps the Schwarzschild case in the early and late times.
@@
Figure 8: The values of $\mathcal{R}_{n}$, $\Omega_{n}$ and $\lambda_{n}$ extracted from the waveform of the damping oscillation observed at $r=60M_{1}$ for $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$, $\mathcal{V}_{2}=100M_{1}$ and $\ell=4$. The yellow and orange solid lines show the predictions of the PL analysis $(2\ell+1)\mathcal{R}_{\rm PL}$ and corresponding Schwarzschild case, respectively. For the bottom panel, since the predicted value from the PL analysis does not match the Schwarzschild case in the early and late times, we multiply the correction factor (blue dashed line) so that it can overlap with the Schwarzschild case in the early and late times. The solid red line shows the corrected PL prediction with the time shifted to that at which the radial null geodesic emanated from the corresponding point on the photon sphere reaches $r=r_{\rm obs}=60M_{1}$.
@@
In the bottom panel of Fig. 8, we can find that the sequence of $\mathcal{R}_{n}$ also has a qualitatively similar behavior to the Penrose limit prediction $\mathcal{R}_{\rm PL}$. However, the significance of the deviation from the constant value is quantitatively about 3 times different from $\mathcal{R}_{\rm PL}$. In the remaining part of this paper, we carefully investigate the parameter dependence of $\mathcal{R}_{n}$, and compare it to $\mathcal{R}_{\rm PL}$.
@@
We calculate the values $\Omega_{\rm ad}$ and $\lambda_{\rm ad}$, and connect them to an effective complex frequency $\omega_{\rm PL,ad}$. Therefore, for $\omega_{\rm PL,ad}$ to be good approximation, the adiabatic approximation must be valid. Since the value $M_{2}=3M_{1}$ in the previous example is likely to be too large, let us consider the case $M_{2}=1.1M_{1}$, $\mathcal{V}_{1}=0$, $\mathcal{V}_{2}=100M_{1}$, and $\ell=4$ as a reference. First, let us check the dependence of the results on the value of $r_{\rm obs}$. In Fig. 9, we show the values $\mathcal{R}_{n}$.
@@
Figure 9: $r_{\rm obs}$ dependence of the values of $\mathcal{R}_{n}$. The value of $\mathcal{R}_{n}$ are plotted for $\mathcal{V}_{\rm ini}=-100M_{1}$, $50M_{1}$ and 0 for each value of $r_{\rm obs}$. The orange line shows the value in the Schwarzschild case. The yellow line shows the corresponding value obtained from the numerical calculation for the static conformal metric of the Vaidya spacetime with the constant accretion rate $M^{\prime}=\pi/2000$. The dashed blue line shows the value given by the Penrose limit analysis with the constant accretion rate $M^{\prime}=\pi/2000$.
@@
Figure 11: The values of $\mathcal{R}_{n}$ extracted from the waveform of the damping oscillation observed at $r=420M_{1}$ for $\mathcal{V}_{1}=0$, $\mathcal{V}_{2}=100M_{1}$, $\ell=4$. The left and right panels are for $M_{2}=1.1M_{1}$ and $M_{2}=1.02M_{1}$, respectively.
We find that the feature quantitatively approaches the prediction from the Penrose limit analysis compared to the case $M_{2}=3M_{1}$. However, the result does not asymptote to the Penrose limit prediction only by changing the value of $M_{2}$. Since the time-domain calculation agrees with the frequency domain and PL analyses with the correction factor $\mathcal{C}$ in the case of constant accretion rate, this disagreement is expected to be caused by the time-dependent accretion rate during the wave propagation between the UCOP and the observer. That is, the difference between the PL limit prediction and the time-domain results is mainly caused by the scattering process during the propagation, and cannot be captured by a local geometry around the UCOP. Then, let us change the period of the accretion, keeping the value of the maximum accretion rate $M^{\prime}_{\rm max}$. The results for the cases $\mathcal{V}_{2}=150M_{1}$ and $\mathcal{V}_{2}=200M_{1}$ with $M^{\prime}_{\rm max}=\pi/2000$ are shown in Fig. 12.
Figure 12: The values of $\mathcal{R}_{n}$ extracted from the waveform of the damping oscillation observed at $r=420M_{1}$ for $\mathcal{V}_{1}=0$, $M^{\prime}_{\rm max}=\pi/2000$, $\ell=4$. The left and right panels are for $\mathcal{V}_{2}=150M_{1}$ and $\mathcal{V}_{2}=200M_{1}$, respectively.
For a sufficiently large value of $\mathcal{V}_{2}$, the values of $\mathcal{R}_{n}$ extracted from the waveform of the damping oscillation seems to approach the prediction from the Penrose limit analysis with the correction factor $\mathcal{C}$. This tendency can be understood as a result of the decrease in the significance of the scattering effects during the wave propagation.
@@
Figure 13: The values of $\mathcal{R}_{n}$ extracted from the waveform of the damping oscillation observed at $r=60M_{1}$ for $\mathcal{V}_{1}=0$, $\mathcal{V}_{2}=100M_{1}$, and $M_{2}=1.1M_{1}$. The left and right panels are for $\ell=6$ and $\ell=10$, respectively.
@@ V. Summary and conclusions
For the static background spacetime, the ringdown waves are known to originate from the damping oscillation modes around the unstable circular orbit of photons (UCOP). This behavior can be justified considering quasinormal modes in the Penrose limit geometry around the null geodesic on the UCOP as is reviewed in Sec. 2. The Penrose limit geometry just provide us the information of the local geometry near the UCOP. Therefore, the scattering effects during the propagation from UCOP to the observer cannot be captured in the Penrose limit analyses. Nevertheless, in the case of static spacetimes, considering the quasinormal mode frequency associated with the static Killing observer, the result can be directly interpreted as the waveform measured by a distant observer. Therefore, the Penrose limit analyses can be also easily compared with the full quasinormal mode analyses.
Our main purpose was to clarify whether this observation also applies to dynamical spacetimes. In this paper, we carefully analyzed the ringdown waveforms in a Vaidya spacetime with time-independent (Sec. 3) and dependent accretion rates (Sec. 4), and compared them with the prediction from the Penrose limit analyses. In the case of a constant accretion rate, the spacetime metric can be described as a conformally static form. Then, both the frequency domain quasinormal mode analysis and the Penrose limit analysis can be performed in the conformal static spacetime. We found that these analyses agree with each other at a certain level when the correction term evaluated by using the Schwarzschild case is taken into account. We have also compared the evaluated quasinormal mode frequencies with those from the time-domain analyses performed in the full Vaidya spacetime with the observer trajectory adjusted to the frequency domain analysis, and found again good agreements. This observation shows us that, even under the existence of the accretion, the Penrose limit analysis may give us the local geometrical origin of the ringdown waveforms.
@@
The temporal frequency monotonically changes from the Schwarzschild value with the initial mass $M_{1}$ to that with the final mass $M_{2}$. This behavior is shared with the Penrose limit analysis, but we observed substantial deviation from the Penrose limit prediction in the time dependence of the temporal frequency. This substantial deviation is mainly caused by the scattering effects during the propagation in a broad sense including the redshift effect. The redshift effect is the dominant effect and equally shifts the real and imaginary parts of the frequency. Since the redshift effect is an additional effect to the original emission process around the photon sphere, the redshift effect causes a deviation between the Penrose limit prediction and the actual waveform. It should be noted that, in a static case, the redshift effect is trivial and automatically taken into account by considering the frequency for the time coordinate associated with the static Killing vector field. Therefore, the redshift effect can be also regarded as a peculiar effect in the dynamical spacetime, and it contains the information about the accretion.
Since the redshift equally affects to the real and imaginary parts of the complex frequency, we may extract the purely geometrical information around the dynamical photon sphere by taking the ratio between the imaginary ($\omega^{\rm Im}$) and real ($\omega^{\rm Re}$) parts of the frequency. We found the characteristic temporal decrease in the ratio $\mathcal{R}=\omega^{\rm Im}/\omega^{\rm Re}$ both in the Penrose limit analysis and the time-domain numerical simulation. It has been shown that, for sufficiently large angular momentum quantum number $\ell$ ($\ell\geq 4$) and observer radius $r_{\rm obs}$ ($r_{\rm obs}\gtrsim 400M_{1}$), the time-dependence of $\mathcal{R}$ approaches to the Penrose limit prediction in the limit $M^{\prime\prime}M_{1}\ll M^{\prime}$ with the prime “ ” being the time derivative. This result indicates that there is still a remaining scattering effect associated with the time-dependence of the accretion rate, and without this remaining scattering effect in an appropriate limit, the Penrose limit prediction may give a good approximation for $\mathcal{R}$. It implies that, in principle, the ringdown waveform can deliver the information of the geometry around the photon sphere to a distant observer even in a dynamical case.

### v3

@@ A. Penrose limit
$\omega^{\rm Re}_{\rm PL,ad}$ $:=$ $\frac{\Omega_{\rm ad}}{2}\left(2\ell+1\right),$ (4.15)
$\omega^{\rm Im}_{\rm PL,ad}$ $:=$ $\frac{\lambda_{\rm ad}}{2},$ (4.16)
$\mathcal{R}^{\rm ad}_{\rm PL}$ $:=$ $\left|\frac{\omega^{\rm Im}_{\rm PL,ad}}{\omega^{\rm Re}_{\rm PL,ad}}\right|.$ (4.17)
@@
Figure 6: The values of $\lambda_{\rm ad}$ and $\Omega_{\rm ad}$ with $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$ and $\mathcal{V}_{2}=100M_{1}$ as a function of $\mathcal{V}$ (left panel) and $\mathcal{V}_{\rm obs}$ (right panel). The null dust is accreting in the shaded region. The ratio $\lambda_{\rm ad}/\Omega_{\rm ad}$ is also plotted with the scale of the right-hand vertical axis as a reference. The red dashed lines in the right panel show all lines in the left panel as references.
We can find that the ratio $\lambda_{\rm ad}/\Omega_{\rm ad}$ is smaller than 1 in the accreting domain. Therefore, if the feature is reflected in the ringdown waveform, we may observe a temporary reduction in the value of $\mathcal{R}_{n}(2\ell+1)$ from the value of the Schwarzschild case. It would be interesting to note that we trivially obtain the Schwarzschild metric, taking the adiabatic approximation for the original Vaidya metric, since we ignore the time-dependence of the mass function. Then we may expect the constant value of the ratio between $\Omega$ and $\lambda$. However, in the adiabatic limit of the PL geometry, the ratio is not constant. That is, the adiabatic limit of the PL geometry reflects more details of the dynamical geometry compared to the naive adiabatic limit of the original geometry.
@@ B. Ringdown in Vaidya spacetimes and comparison with the Penrose limit analysis
We plot the values of $\Omega_{n}:=2\omega^{\rm Re}_{n}/(2\ell+1)$, $\lambda_{n}:=2\omega^{\rm Im}_{n}$ and $\mathcal{R}_{n}$, with $\mathcal{V}_{\rm obs}(\mathcal{V}_{n})$ being the horizontal axis value in Fig. 8. As in the case of Fig. 3, the amplitude of the damping oscillation drops quickly, so that the numerical plots in the late time scatter a lot (see red dots in Fig. 8) due to the numerical error. Therefore, we plot the results with several different values of the initial time $\mathcal{V}_{\rm ini}$, and focus on the time-domain in which the waveform exhibits an idealized exponentially damping oscillation. In the top panels, we can find that the values of $\Omega_{n}$ and $\lambda_{n}$ significantly deviate from $\Omega_{\rm ad}$ and $\lambda_{\rm ad}$, respectively. Since the qualitative behaviors are similar in both $\Omega_{n}$ and $\lambda_{n}$, these deviations mainly originate from the redshift effect, which equally causes the time delay to the real and imaginary parts of the frequency during the propagation from the UCOP to the observer. To extract the effects independently of the redshift, let us focus on the ratio $\mathcal{R}_{n}$ between the imaginary and real frequencies shown in the bottom panel of Fig. 8. For the bottom panel, the predicted value from the PL analysis does not match the Schwarzschild case in both of the early and late times because of the deviation presented in Table 2. Thus, we multiply the correction factor $\mathcal{C}$ (blue dashed line), so that it overlaps the Schwarzschild case in the early and late times.
@@
Figure 8: The values of $\mathcal{R}_{n}$, $\Omega_{n}$ and $\lambda_{n}$ extracted from the waveform of the damping oscillation observed at $r=60M_{1}$ for $M_{2}=3M_{1}$, $\mathcal{V}_{1}=0$, $\mathcal{V}_{2}=100M_{1}$ and $\ell=4$. In each panel, different kinds of dots show the results obtained from the numerical simulation for different values of $\mathcal{V}_{\rm ini}$ indicated in the legend. In the top panels, the yellow lines show the values of $\Omega_{\rm ad}$ and $\lambda_{\rm ad}$, respectively, obtained by the PL analysis. The orange solid lines show the values of $\Omega_{\rm Sch}M_{1}/M(\mathcal{V})$ and $\lambda_{\rm Sch}M_{1}/M(\mathcal{V})$ with $\Omega_{\rm Sch}=2\omega_{\rm Sch}^{\rm Re}/(2\ell+1)$ and $\lambda_{\rm Sch}=2\omega_{\rm Sch}^{\rm Im}$. In the bottom panel, the yellow dashed and orange solid lines show the predictions of the PL analysis $(2\ell+1)\mathcal{R}_{\rm PL}$ and corresponding Schwarzschild case, respectively. For the bottom panel, since the predicted value from the PL analysis does not match the Schwarzschild case in the early and late times, we multiply the correction factor (blue dashed line) so that it can overlap with the Schwarzschild case in the early and late times. The solid red line shows the corrected PL prediction with the time shifted to that at which the radial null geodesic emanated from the corresponding point on the photon sphere reaches $r=r_{\rm obs}=60M_{1}$.
@@
In the bottom panel of Fig. 8, we can find that the sequence of $\mathcal{R}_{n}$ also has a qualitatively similar behavior to the PL prediction $\mathcal{R}_{\rm PL}$. However, the significance of the deviation from the constant value is quantitatively about 3 times different from $\mathcal{R}_{\rm PL}$. In the remaining part of this paper, we carefully investigate the parameter dependence of $\mathcal{R}_{n}$, and compare it to $\mathcal{R}_{\rm PL}$.
@@
We calculate the values $\Omega_{\rm ad}$ and $\lambda_{\rm ad}$, and connect them to an effective complex frequency $\omega_{\rm PL,ad}$. Therefore, for $\omega_{\rm PL,ad}$ to be a good approximation, the adiabatic approximation must be valid. Since the value $M_{2}=3M_{1}$ in the previous example is likely to be too large, let us consider the case $M_{2}=1.1M_{1}$, $\mathcal{V}_{1}=0$, $\mathcal{V}_{2}=100M_{1}$, and $\ell=4$ as a reference. First, let us check the dependence of the results on the value of $r_{\rm obs}$. In Fig. 9, we show the values $\mathcal{R}_{n}$.
Figure 9: $r_{\rm obs}$ dependence of the values of $\mathcal{R}_{n}$. The values of $\mathcal{R}_{n}$ are plotted for $\mathcal{V}_{\rm ini}=-100M_{1}$ (blue), $50M_{1}$ (green) and 0 (purple)
@@
for each value of $r_{\rm obs}$. The orange line shows the value in the Schwarzschild case. The yellow line shows the corresponding value obtained from the numerical calculation for the static conformal metric of the Vaidya spacetime with the constant accretion rate $M^{\prime}=\pi/2000$. The dashed blue line shows the value given by the PL analysis with the constant accretion rate $M^{\prime}=\pi/2000$.
@@
Figure 11: The values of $\mathcal{R}_{n}$ extracted from the waveform of the damping oscillation observed at $r=420M_{1}$ for $\mathcal{V}_{1}=0$, $\mathcal{V}_{2}=100M_{1}$, $\ell=4$. The values of $\mathcal{R}_{n}$ are plotted for $\mathcal{V}_{\rm ini}=-100M_{1}$ (blue), $50M_{1}$ (green) and 0 (purple). The left and right panels are for $M_{2}=1.1M_{1}$ and $M_{2}=1.02M_{1}$, respectively.
We find that the feature quantitatively approaches the prediction from the PL analysis compared to the case $M_{2}=3M_{1}$. However, the result does not asymptote to the PL prediction only by changing the value of $M_{2}$. Since the time-domain calculation agrees with the frequency domain and PL analyses with the correction factor $\mathcal{C}$ in the case of constant accretion rate, this disagreement is expected to be caused by the time-dependent accretion rate during the wave propagation between the UCOP and the observer. That is, the difference between the PL limit prediction and the time-domain results is mainly caused by the scattering process during the propagation, and cannot be captured by a local geometry around the UCOP. Then, let us change the period of the accretion, keeping the value of the maximum accretion rate $M^{\prime}_{\rm max}$. The results for the cases $\mathcal{V}_{2}=150M_{1}$ and $\mathcal{V}_{2}=200M_{1}$ with $M^{\prime}_{\rm max}=\pi/2000$ are shown in Fig. 12.
Figure 12: The values of $\mathcal{R}_{n}$ extracted from the waveform of the damping oscillation observed at $r=420M_{1}$ for $\mathcal{V}_{1}=0$, $M^{\prime}_{\rm max}=\pi/2000$, $\ell=4$. The values of $\mathcal{R}_{n}$ are plotted for $\mathcal{V}_{\rm ini}=-100M_{1}$ (blue), $50M_{1}$ (green) and 0 (purple). The left and right panels are for $\mathcal{V}_{2}=150M_{1}$ and $\mathcal{V}_{2}=200M_{1}$, respectively.
For a sufficiently large value of $\mathcal{V}_{2}$, the values of $\mathcal{R}_{n}$ extracted from the waveform of the damping oscillation seems to approach the prediction from the PL analysis with the correction factor $\mathcal{C}$. Once we accept the existence of the limit at which the results asymptote to the PL prediction, we may expect that the deviation from the PL prediction would be caused by some kind of scattering effect during the wave propagation. Then, our results suggest that the significance of this scattering effect decreases for a longer period of the accretion with a fixed value of $M^{\prime}_{\rm max}$.
@@
Figure 13: The values of $\mathcal{R}_{n}$ extracted from the waveform of the damping oscillation observed at $r=60M_{1}$ for $\mathcal{V}_{1}=0$, $\mathcal{V}_{2}=100M_{1}$, and $M_{2}=1.1M_{1}$. The values of $\mathcal{R}_{n}$ are plotted for $\mathcal{V}_{\rm ini}=-100M_{1}$ (blue), $50M_{1}$ (green) and 0 (purple). The left and right panels are for $\ell=6$ and $\ell=10$, respectively.
Our results indicate that the value $\ell=4$ is sufficiently large so that the results will almost converge to the high-frequency limit. Thus, in the dynamical case shown here, for the realization of the PL limit, the large $\ell$ limit is not a sufficient condition, but we need additional conditions for the accretion rate and period.
@@ V. Summary and conclusions
For the stationary background spacetime, the ringdown waves are known to originate from the damping oscillation modes around the UCOP. This behavior can be justified considering QNMs in the PL geometry around the null geodesic on the UCOP, as is reviewed in Sec. 2. The PL geometry just provide us the information of the local geometry near the UCOP. Therefore, the scattering effects during the propagation from UCOP to the observer cannot be captured in the PL analyses. Nevertheless, in the case of stationary spacetimes, considering the QNM frequency associated with the stationary Killing observer, the result can be directly interpreted as the waveform measured by a distant observer. Therefore, the PL analyses can be also easily compared with the full QNM analyses.
Our main purpose was to clarify whether this observation also applies to dynamical spacetimes. In this paper, we carefully analyzed the ringdown waveforms in a Vaidya spacetime with time-independent (Sec. 3) and dependent accretion rates (Sec. 4), and compared them with the prediction from the PL analyses. In the case of a constant accretion rate, the spacetime metric can be described as a conformally static form. Then, both the frequency domain QNM analysis and the PL analysis can be performed in the conformal static spacetime. We found that these analyses agree with each other at a certain level when the correction term evaluated by using the Schwarzschild case is taken into account. We have also compared the evaluated QNM frequencies with those from the time-domain analyses performed in the full Vaidya spacetime with the observer trajectory adjusted to the frequency domain analysis, and found good agreement again. This observation shows us that, even under the existence of the accretion, the PL analysis may give us the local geometrical origin of the ringdown waveforms.
@@
The temporal frequency monotonically changes from the Schwarzschild value with the initial mass $M_{1}$ to that with the final mass $M_{2}$. This behavior is shared with the PL analysis, but we observed substantial deviation from the PL prediction in the time dependence of the temporal frequency. This substantial deviation is mainly caused by the scattering effects during the propagation in a broad sense including the redshift effect. The redshift effect is the dominant effect and equally shifts the real and imaginary parts of the frequency. Since the redshift effect is an additional effect to the original emission process around the photon sphere, the redshift effect causes a deviation between the PL prediction and the actual waveform. It should be noted that, in a stationary case, the redshift effect is trivial and automatically taken into account by considering the frequency for the time coordinate associated with the stationary Killing vector field. Therefore, the redshift effect can be also regarded as a peculiar effect in the dynamical spacetime, and it contains the information about the accretion.
Since the redshift equally affects to the real and imaginary parts of the complex frequency, we may extract the purely geometrical information around the dynamical photon sphere by taking the ratio between the imaginary ($\omega^{\rm Im}$) and real ($\omega^{\rm Re}$) parts of the frequency. We found the characteristic temporal decrease in the ratio $\mathcal{R}=\omega^{\rm Im}/\omega^{\rm Re}$ both in the PL analysis and the time-domain numerical simulation. It has been shown that, for sufficiently large angular momentum quantum number $\ell$ ($\ell\geq 4$) and observer radius $r_{\rm obs}$ ($r_{\rm obs}\gtrsim 400M_{1}$), the time-dependence of $\mathcal{R}$ approaches to the PL prediction under an additional condition $M^{\prime\prime}M_{1}\ll M^{\prime}$ with the prime “ ” being the time derivative. This result indicates that there is still a remaining scattering effect associated with the time-dependence of the accretion rate, and without this remaining scattering effect in an appropriate limit, the PL prediction may give a good approximation for $\mathcal{R}$. It implies that, in principle, the ringdown waveform can deliver the information of the geometry around the photon sphere to a distant observer even in a dynamical case.